Grade 12Mechanics

Two friends starts from same point with constant acceleration of 5 metre per second square second one starts journey 5 seconds later find the time taken by the first one such that he is hundred metre of second one

Jyothirmai Kakarla
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
If the time travelled upto the given point by second one istseconds, the first one has takent+5seconds.

Since the difference between their distances travelled is 100m,

\dfrac{1}{2}a(t+5)^2-\dfrac{1}{2}at^2=100
\implies \dfrac{1}{2}a(10t+25)=100
\implies t=1.5s
Hence time taken by first one is6.5s